OpenStudy (anonymous):

Write the slope-intercept equation for the line with slope -2/5 and y-intercept (0, -3) and graph the line

OpenStudy (anonymous):

( y - -3) = (-2/5) ( x - 0 ) y + 3 = -2/5x y = 2/5x - 3

OpenStudy (anonymous):

how would i graph that

OpenStudy (anonymous):

Your y-intercept = -3 So make a point at (0, -3) From there, use your slope -2/5 So from your point, you go down 2 and over to the right 5. Or you could go up 2 and go over to the left 5